# Martin’s Conjecture in the Enumeration Degrees

### v1

@@ Abstract.
Martin’s Conjecture is that every definable function on the Turing degrees is either constant or increasing, and that every increasing function is an iterate of the Turing jump. This classification has already been corroborated for the class of uniformly invariant functions and a long-standing conjecture by Steel is that every definable function on the Turing degrees is equivalent to a uniformly invariant one. We explore whether a similar classification is possible in the enumeration degrees, an extension of the Turing degrees. We show that the spectrum of behaviours is much wider in the enumeration degrees, even for uniformly invariant functions. However, our main result is that uniformly invariant functions behave locally as nicely as possible: they are constant, increasing, or above the skip operator. As a consequence, we show that there is a definable function in the enumeration degrees that is not equivalent to a uniformly invariant one.
@@ 1. Introduction
This theorem gives us a way to compare the asymptotic behaviour of functions from $\mathcal{P}(\mathbb{N})$ to $\mathcal{P}(\mathbb{N})$ relative to Turing reducibility. Namely, we say that $f$ is Turing-below $g$ on a cone, denoted by $f\leq_{T}^{\triangledown}g$, if there is some $A\subseteq\mathbb{N}$ such that $f(X)\leq_{T}g(X)$ for all $X\geq_{T}A$. Similarly, $f$ is increasing on a cone if $\operatorname{Id}\leq^{\triangledown}_{T}f$ and constant on a cone if $f(Z)\equiv_{T}f(Y)$, for all $X$ and $Y$ in some cone.
@@
We are interested in Martin’s Conjecture in the context of the enumeration degrees due to their close structural relationship to the Turing degrees. A set $A$ is enumeration reducible to $B$, denoted by $A\leq_{e}B$, if every enumeration of $B$ computes an enumeration of $A$. The structure of the enumeration degrees $\mathcal{D}_{e}$ is then defined in the usual way as the set of equivalence classes of the symmetric closure of enumeration reducibility. The Turing degrees $\mathcal{D}_{T}$ embed into the enumeration degrees [23]; moreover, this copy of the Turing degrees is first-order definable in $\mathcal{D}_{e}$ (with the language $\{\leq\}$) [6]. So, the $e$-invariant functions provide a broader playground to test the uniqueness of the Turing jump. However, in the enumeration degrees there is no single candidate to take the place of the Turing jump. In fact, there are two well-studied uniformly $e$-invariant extensions of the Turing jump to the enumeration degrees: the enumeration jump—originally defined by Cooper [8]—and the enumeration skip, introduced by Andrews et al. [2]. While both operators share common properties with the Turing jump, each of them presents a distinct global behaviour. For instance, the enumeration jump is increasing while the skip is not, and the skip satisfies Friedberg’s jump inversion but the enumeration jump does not. Given this state of affairs, we explore the extent to which the uniqueness of the Turing jump is inherited by the enumeration jump and the skip, and whether the phenomenon conjectured by Martin is reflected in the enumeration degrees.
@@
Section 4 focuses on uniformly $e$-invariant functions and their uniformity functions. The construction from Section 3 produces uniformly invariant functions only because they are locally constant (i.e., the image of every enumeration degree is a single set). Note that in the Turing degrees, AD implies that locally constant functions are constant in some cone. However, through an analysis of uniformity functions, we show that any countable family of uniformly $e$-invariant functions can be merged under a single uniformity function. As a consequence, the preorder $\leq_{e}^{\triangledown}$ has infinite antichains of uniformly $e$-invariant functions, even if we ignore the locally constant functions.
@@
As a consequence, we extend Theorem 1.3 to uniformly invariant functions from the Turing degrees to the enumeration degrees. Theorem 5.1 also provides insight on the difference between invariance and uniform invariance. In Section 6, we produce a Borel $e$-invariant function that is not uniformly $e$-invariant on any upper cone by taking advantage of another structural difference between the Turing degrees and the enumeration degrees: the existence of $\mathcal{K}$-pairs. The lack of uniformity comes from an incompatibility between the combinatorics of $\mathcal{K}$-pairs and Theorem 5.1. As far as we know, this is the first such example in any degree structure.
@@ 2. Enumeration Degrees and the Skip
$x\in A\quad\mbox{if and only if}\quad\exists u(\langle n,u\rangle\in\Gamma\;\&\;D_{u}\subseteq B)\,$,
where $\{D_{n}\}$ is the standard numbering of finite sets. In this context, we say that $\Gamma$ is an enumeration operator (or just an $e$-operator) and we write $A=\Gamma(B)$. To simplify notation, we identify a finite set $D$ with its canonical index $u$. So, we write $\langle n,D\rangle\in\Gamma$ instead of $\langle n,u\rangle\in\Gamma$.
@@
As usual, one can define enumeration equivalence by setting $A\equiv_{e}B$ when $A\leq_{e}B$ and $B\leq_{e}A$. If $A\leq_{e}B$ via $i$ and $B\leq_{e}A$ via $j$, we say that $A\equiv B$ via $\langle i,j\rangle$. Enumeration equivalence is an equivalence relation and induces the degree structure $\mathcal{D}_{e}=\mathcal{P}(\mathbb{N})/\equiv_{e}$ which we call the enumeration degrees. It is an upper semi-lattice with least element $\bm{0}_{e}$, the degree of c.e. sets, and least upper bound given by the usual join operator.
@@
Let $A,G\subseteq\mathbb{N}$ such that $G$ is $\langle A\rangle$-generic. Then $G\oplus A$ is a strong quasiminimal cover.
@@
There is another class of enumeration degrees that will be of particular interest to us—the cototal degrees. They were studied by Andrews et. al. [2] where they showed a broad range of applications to effective mathematics. A set $A$ is cototal if $A\leq_{e}\overline{A}$ and an enumeration degree is cototal if it contains a cototal set. Every total degree $\bm{a}$ is cototal because if $A\in\bm{a}$ is total, then $A\equiv_{e}A\oplus\overline{A}\equiv_{e}\overline{A}\oplus A=\overline{A\oplus\overline{A}}$. However, there are cototal degrees that are not total. Gutteridge [14] and Sorbi [26] independently constructed cototal quasiminimal degrees (a degree is quasiminimal if it is a quasiminimal cover of $\bm{0}_{e}$, the enumeration degree of c.e. sets). These constructions can be relativized to total degrees, showing that the cototal degrees that are not total are also cofinal. Moreover, most sets are not cototal, both in terms of measure and category. Li [19] showed that the weakly 3-random sets cannot have cototal degree, concluding that the cototal degrees have measure zero, and Andrews et. al. [2] proved that the enumeration degrees of 2-generic sets are not cototal, which implies that they are meager.
@@
An alternative extension of the Turing jump is given by the skip operator. This operator was only recently studied by Andrews et. al. [2], and while it also shares many useful properties with the Turing jump, it presents quite a distinct behaviour. The skip of $A$ is defined as the set $A^{\diamond}=\overline{K}_{A}$ and similarly to the enumeration jump, it is a uniformly $e$-invariant function. The degree of $A^{\diamond}$ is the maximum degree of a set $\overline{B}$ such that $B\equiv_{e}A$. From this perspective, it is not surprising that cototality and the skip operator are closely related.
Proposition 2.4 (Andrews et. al. [2]).
@@ 3. Global Failure
Let $\mathcal{C}$ be the class of all sets with cototal degree and define $h:\mathcal{P}(\mathbb{N})\to\mathcal{P}(\mathbb{N})$} as
@@
There is an arithmetical function
$g\colon\mathcal{P}(\mathbb{N})\times\mathcal{P}(\mathbb{N})\to\mathcal{P}(\mathbb{N})$} such that
@@
Note that to satisfy a $\mathcal{Q}$ requirement for every $X\in\deg(A)$ at the same time, we will add infinitely many elements to $G$, but only finitely many at each column. There is, however, a problem that arises from attempting to carry out the construction simultaneously for all $X\in\deg_{T}(A)$. If $X\equiv_{T}Z$ via $\langle j,i\rangle$, both the requirements $\mathcal{P}^{Z}_{e,i,j}$ and $\mathcal{P}^{X}_{e,i,j}$ might try to modify $G_{X}$ simultaneously in incompatible ways. More generally, a conflict may arise between the $\mathcal{P}^{X}_{e,i,j}$ strategy and the $\mathcal{P}^{Z}_{e,i,j}$ strategy when $Z$ is obtained from a repeated application of $\Phi_{j}$ to $X$, or when $X$ is obtained from a repeated application of $\Phi_{i}$ to $Z$. We solve this problem by establishing, for each pair $\langle i,j\rangle$, a well-founded partial order $\prec_{i,j}$ on $\deg_{T}(A)$ that ranks the requirements by priority. Namely, $\mathcal{P}^{X}_{e,i,j}$ has higher priority than $\mathcal{P}^{Z}_{e,i,j}$ if $X\prec_{i,j}Z$. While $\prec_{i,j}$ will not be a linear order, it will resolve the situations when a conflict could happen.
@@
To see that $f$ is Borel, observe that for $X\subseteq\mathbb{N}$, we can decide if there is some $A$ such that $\deg_{e}(X)\in\mathcal{C}(A)$ by first checking if $\deg_{e}(X)$ is a strong quasiminimal cover of some total degree $\bm{a}$ and, in case it is, verifying if $X\equiv_{e}M_{Y}$ for some $Y\oplus\overline{Y}\in\bm{a}$. Checking if $\deg_{e}(X)$ is a strong quasiminimal cover of some total degree $\bm{a}$ is $\Sigma^{0}_{5}(X)$: we need to check that $X$ does not have total degree—which is $\Sigma^{0}_{3}(X)$—and find if there is some total set $Y\oplus\bar{Y}\leq_{e}X$ such that for any enumeration operator $\Gamma$ with $\Gamma(X)$ total, we have that $\Gamma(X)\leq_{e}Y\oplus\bar{Y}$. Writing down the statement we get a natural $\Sigma^{0}_{5}$ definition:
@@
If $X$ is not a strong quasiminimal cover, then $f(X)=\varnothing$. Otherwise, we have to check if $X=M_{Y}$ for some set $Y\oplus\bar{Y}$ in the degree that $X$ is a strong quasiminimal cover of. In the proof of Theorem 3.2, we saw that $Y^{(4)}$ computes $\bigoplus_{Z\in\deg_{T}(Y)}M_{Z}$, so $X^{(4)}$ also computes it. Thus, $X^{(5)}$ can answer if there is a $Y$ such that $X=M_{Y}$ and, if the answer is “yes”, compute $A=\{n\mid 2n\in Y\}$ such that $X\in\mathcal{C}(A)$; in other words, $X^{(5)}$ can decide in which case of the definition of $f$ we are, so $f$ is Borel. ∎
@@
is uniformly $e$-invariant and $g_{A}\leq_{e}^{\triangledown}f$
@@ 4. Uniform Invariance
Let $f$ be uniformly $e$-invariant with uniformity function $u$. If $A\equiv_{e}B$ via $\langle e,i\rangle$ and $B\equiv C$ via $\langle j,k\rangle$, then $f(A)\equiv_{e}f(C)$ via $u(e,i)\circ u(j,k)$.
@@
Let $\mathcal{C}$ be the class of all sets with cototal degree and define $h:\mathcal{P}(\mathbb{N})\to\mathcal{P}(\mathbb{N})$ as
@@ 5. A Positive Local Result
We have already seen that the global structure of the $e$-invariant functions is not as tame as Martin conjectured for the $T$-invariant case, even with the uniformity assumption. However, the failure comes from a topological accident; we can assemble together different functions to create many unique patchworks on every upper cone. The empirical observation remains: the only known uniformly $e$-invariant functions are combinations of constants, the identity, and iterates of the skip or the jump. One approach that isolates this problem is to follow Bard’s proof [4] of part 1 of the uniform Martin’s Conjecture (UMC 1). Adapting his approach to the context of the enumeration degrees, we obtain the best possible local analogue of (UMC 1) in the enumeration degrees (Theorem 5.1). So, despite the negative results in the previous sections, the enumeration jump and the skip might indeed generate all natural uniformly $e$-invariant functions that are not locally constant, even if the generating process is not just iteration, we will discuss this further in the last section.
@@
If $f$ is uniformly $e$-invariant and $f(A)<_{e}A$, then $f\restriction\deg_{e}(A)$ is constant. Thus, the only regressive uniformly $e$-invariant functions are locally constant.
@@
Now, let $\mathcal{T}=\{A\oplus\overline{A}\mid A\subseteq\mathbb{N}\}$. The function $f$ induces a uniformly $e$-invariant function $\hat{f}:\mathcal{T}\to\mathcal{P}(\mathbb{N})$ given by $\hat{f}(A\oplus\bar{A})=f(A)$. Moreover, we have that $\hat{f}\restriction\deg_{e}(A\oplus\overline{A})$ is not constant for any $A\oplus\overline{A}\geq_{e}Z\oplus\overline{Z}$. It is straightforward to modify the proof of Theorem 5.1 to accept functions only defined on sets of the form $A\oplus\overline{A}$. Thus, $\hat{f}(A\oplus\overline{A})\geq_{e}A\oplus\overline{A}$ or $\hat{f}(A\oplus\overline{A})\geq_{e}(A\oplus\overline{A})^{\diamond}\equiv_{e}(A\oplus\overline{A})^{\prime}$. In any case, $A\leq_{e}A\oplus\overline{A}\leq_{e}\hat{f}(A\oplus\overline{A})=f(A)$. ∎

### v2

@@ Abstract.
Martin’s Conjecture states that every definable function on the Turing degrees is either constant or increasing, and that every increasing function is an iterate of the Turing jump. This classification has already been corroborated for the class of uniformly invariant functions and a long-standing conjecture by Steel is that every definable function on the Turing degrees is equivalent to a uniformly invariant one. We explore whether a similar classification is possible in the enumeration degrees, an extension of the Turing degrees. We show that the spectrum of behaviour is much wider in the enumeration degrees, even for uniformly invariant functions. However, our main result is that uniformly invariant functions behave locally as nicely as possible: they are constant, increasing, or above the skip operator. As a consequence, we show that there is a definable function in the enumeration degrees that is not equivalent to a uniformly invariant one on any cone.
@@ 1. Introduction
This theorem gives us a way to compare the asymptotic behaviour of functions from $\mathcal{{P}}(\mathbb{{N}})$ to $\mathcal{{P}}(\mathbb{{N}})$ relative to Turing reducibility. Namely, we say that $f$ is Turing-below $g$ on a cone, denoted by $f\leq_{T}^{\triangledown}g$, if there is some $A\subseteq\mathbb{{N}}$ such that $f(X)\leq_{T}g(X)$ for all $X\geq_{T}A$. Similarly, $f$ is increasing on a cone if $\operatorname{Id}\leq^{\triangledown}_{T}f$ and constant on a cone if $f(X)\equiv_{T}f(Y)$, for all $X$ and $Y$ in some cone.
@@
We are interested in Martin’s Conjecture in the context of the enumeration degrees due to their close structural relationship to the Turing degrees. A set $A$ is enumeration reducible to $B$, denoted by $A\leq_{e}B$, if every enumeration of $B$ computes an enumeration of $A$ (see Definition 2.1). The structure of the enumeration degrees $\mathcal{D}_{e}$ is then defined in the usual way as the set of equivalence classes of the symmetric closure of enumeration reducibility. The Turing degrees $\mathcal{{D}}_{T}$ embed into the enumeration degrees [23]; moreover, this copy of the Turing degrees is first-order definable in $\mathcal{{D}}_{e}$ (with the language $\{\leq\}$) [6]. So, the $e$-invariant functions provide a broader playground to test the uniqueness of the Turing jump. However, in the enumeration degrees there is no single candidate to take the place of the Turing jump. In fact, there are two well-studied uniformly $e$-invariant extensions of the Turing jump to the enumeration degrees: the enumeration jump—originally defined by Cooper [8]—and the enumeration skip, introduced by Andrews et al. [2]. While both operators share common properties with the Turing jump, each of them presents a distinct global behaviour. For instance, the enumeration jump is increasing while the skip is not, and the skip satisfies Friedberg’s jump inversion but the enumeration jump does not. Given this state of affairs, we explore the extent to which the uniqueness of the Turing jump is inherited by the enumeration jump and the skip, and whether the phenomenon conjectured by Martin is reflected in the enumeration degrees.
@@
Section 4 focuses on uniformly $e$-invariant functions and their uniformity functions. The construction from Section 3 produces uniformly invariant functions only because they are locally constant; that is, the image of every enumeration degree is a single set. Note that in the Turing degrees, TD implies that locally constant functions are constant in some cone. However, through an analysis of uniformity functions, we show that any countable family of uniformly $e$-invariant functions can be merged under a single uniformity function. As a consequence, the preorder $\leq_{e}^{\triangledown}$ has infinite antichains of uniformly $e$-invariant functions, even if we ignore the locally constant functions.
@@
As a consequence, we extend part 1 of Theorem 1.3 (UMC 1) to uniformly invariant functions from the Turing degrees to the enumeration degrees. Theorem 5.1 also provides insight on the difference between invariance and uniform invariance. In Section 6, we produce a Borel $e$-invariant function that is not equivalent to a Borel uniformly $e$-invariant function on any upper cone by taking advantage of another structural difference between the Turing degrees and the enumeration degrees: the existence of $\mathcal{{K}}$-pairs. The lack of uniformity comes from an incompatibility between the combinatorics of $\mathcal{{K}}$-pairs and Theorem 5.1. As far as we know, this is the first such example in any degree structure.
@@ 2. Enumeration Degrees and the Skip
$n\in A\quad\mbox{if and only if}\quad\exists u(\langle n,u\rangle\in\Gamma\;\&\;D_{u}\subseteq B)\,$,
where $\{D_{u}\}$ is the standard numbering of finite sets. In this context, we say that $\Gamma$ is an enumeration operator (or just an $e$-operator) and we write $A=\Gamma(B)$. To simplify notation, we identify a finite set $D$ with its canonical index $u$. So, we write $\langle n,D\rangle\in\Gamma$ instead of $\langle n,u\rangle\in\Gamma$.
@@
As usual, one can define enumeration equivalence by setting $A\equiv_{e}B$ when $A\leq_{e}B$ and $B\leq_{e}A$. If $A\leq_{e}B$ via $i$ and $B\leq_{e}A$ via $j$, we say that $A\equiv_{e}B$ via $\langle i,j\rangle$. Enumeration equivalence is an equivalence relation and induces the degree structure $\mathcal{{D}}_{e}=\mathcal{{P}}(\mathbb{{N}})/\equiv_{e}$ which we call the enumeration degrees. It is an upper semi-lattice with least element $\bm{0}_{e}$, the degree of c.e. sets, and least upper bound given by the usual join operator.
@@
Let $A,G\subseteq\mathbb{{N}}$ such that $G$ is $\langle A\rangle$-generic. Then $G\oplus A$ is a strong quasiminimal cover of $A$.
@@
There is another class of enumeration degrees that will be of particular interest to us—the cototal degrees. They were studied by Andrews et al. [2] where they showed a broad range of applications to effective mathematics. A set $A$ is cototal if $A\leq_{e}\overline{A}$ and an enumeration degree is cototal if it contains a cototal set. Every total degree $\bm{a}$ is cototal because if $A\in\bm{a}$ is total, then $A\equiv_{e}A\oplus\overline{A}\equiv_{e}\overline{A}\oplus A=\overline{A\oplus\overline{A}}$. However, there are cototal degrees that are not total. Gutteridge [14] and Sorbi [26] independently constructed cototal quasiminimal degrees (a degree is quasiminimal if it is a quasiminimal cover of $\bm{0}_{e}$, the enumeration degree of c.e. sets). These constructions can be relativized to total degrees, showing that the cototal degrees that are not total are also cofinal. Moreover, most sets are not cototal, both in terms of measure and category. Li [19] showed that the weakly 3-random sets cannot have cototal degree, concluding that the cototal degrees have measure zero, and Andrews et al. [2] proved that the enumeration degrees of 2-generic sets are not cototal, which implies that they are meager.
@@
An alternative extension of the Turing jump is given by the skip operator. This operator was only recently studied by Andrews et al. [2], and while it also shares many useful properties with the Turing jump, it presents quite a distinct behaviour. The skip of $A$ is defined as the set $A^{\diamond}=\overline{K}_{A}$ and similarly to the enumeration jump, it is a uniformly $e$-invariant function. The degree of $A^{\diamond}$ is the maximum degree of a set $\overline{B}$ such that $B\equiv_{e}A$. From this perspective, it is not surprising that cototality and the skip operator are closely related.
Proposition 2.4 (Andrews et al. [2]).
@@ 3. Global Failure
Define $h:\mathcal{{P}}(\mathbb{{N}})\to\mathcal{{P}}(\mathbb{{N}})$} as
@@
There is an arithmetical function $g\colon\mathcal{{P}}(\mathbb{{N}})\times\mathcal{{P}}(\mathbb{{N}})\to\mathcal{{P}}(\mathbb{{N}})$} such that
@@
Note that to satisfy a $\mathcal{{Q}}$ requirement for every $X\in\deg_{T}(A)$ at the same time, we will add infinitely many elements to $G$, but only finitely many at each column. There is, however, a problem that arises from attempting to carry out the construction simultaneously for all $X\in\deg_{T}(A)$. If $X\equiv_{T}Z$ via $\langle j,i\rangle$, both the requirements $\mathcal{{P}}^{Z}_{e,i,j}$ and $\mathcal{{P}}^{X}_{e,i,j}$ might try to modify $G_{X}$ simultaneously in incompatible ways. More generally, a conflict may arise between the $\mathcal{{P}}^{X}_{e,i,j}$ strategy and the $\mathcal{{P}}^{Z}_{e,i,j}$ strategy when $Z$ is obtained from a repeated application of $\Phi_{j}$ to $X$, or when $X$ is obtained from a repeated application of $\Phi_{i}$ to $Z$. We solve this problem by establishing, for each pair $\langle i,j\rangle$, a well-founded partial order $\prec_{i,j}$ on $\deg_{T}(A)$ that ranks the requirements by priority. Namely, $\mathcal{{P}}^{X}_{e,i,j}$ has higher priority than $\mathcal{{P}}^{Z}_{e,i,j}$ if $X\prec_{i,j}Z$. While $\prec_{i,j}$ will not be a linear order, it will resolve the situations when a conflict could happen.
@@
To see that $f$ is Borel, observe that for $X\subseteq\mathbb{{N}}$, we can decide if there is some $A$ such that $\deg_{e}(X)\in\mathcal{{C}}(A)$ by first checking if $\deg_{e}(X)$ is a strong quasiminimal cover of some total degree $\bm{a}$ and, in case it is, verifying if $X\equiv_{e}M_{Y}$ for some $Y\oplus\overline{Y}\in\bm{a}$. Checking if $\deg_{e}(X)$ is a strong quasiminimal cover of some total degree $\bm{a}$ is $\Sigma^{0}_{5}(X)$: we need to check that $X$ does not have total degree—which is $\Sigma^{0}_{3}(X)$—and find if there is some total set $Y\oplus\overline{Y}\leq_{e}X$ such that for any enumeration operator $\Gamma$ with $\Gamma(X)$ total, we have that $\Gamma(X)\leq_{e}Y\oplus\overline{Y}$. Writing down the statement we get a natural $\Sigma^{0}_{5}$ definition:
@@
If $X$ is not a strong quasiminimal cover, then $f(X)=\varnothing$. Otherwise, we have to check if $X=M_{Y}$ for some set $Y\oplus\overline{Y}$ in the degree that $X$ is a strong quasiminimal cover of. In the proof of Theorem 3.2, we saw that $Y^{(4)}$ computes $\bigoplus_{Z\in\deg_{T}(Y)}M_{Z}$, so $X^{(4)}$ also computes it. Thus, $X^{(5)}$ can answer if there is a $Y$ such that $X=M_{Y}$ and, if the answer is “yes”, compute $A=\{n\mid 2n\in Y\}$ such that $X\in\mathcal{{C}}(A)$; in other words, $X^{(5)}$ can decide in which case of the definition of $f$ we are, so $f$ is Borel. ∎
@@
is uniformly $e$-invariant and $g_{A}\leq_{e}^{\triangledown}f$. ∎
@@ 4. Uniform Invariance
Let $f$ be uniformly $e$-invariant with uniformity function $u$. If $A\equiv_{e}B$ via $\langle e,i\rangle$ and $B\equiv_{e}C$ via $\langle j,k\rangle$, then $f(A)\equiv_{e}f(C)$ via $u(e,i)\circ u(j,k)$.
@@
Define $h:\mathcal{{P}}(\mathbb{{N}})\to\mathcal{{P}}(\mathbb{{N}})$ as
@@ 5. A Positive Local Result
We have already seen that the global structure of the $e$-invariant functions is not as tame as Martin conjectured for the $T$-invariant case, even with the uniformity assumption. However, the failure comes from a topological accident; we can assemble together different functions to create many unique patchworks on every upper cone. The empirical observation remains: the only known uniformly $e$-invariant functions are combinations of constants, the identity, and iterates of the skip or the jump. One approach that isolates this problem is to follow Bard’s proof [4] of part 1 of the uniform Martin’s Conjecture (UMC 1). Adapting his approach to the context of the enumeration degrees, we obtain the best possible local analogue of (UMC 1) in the enumeration degrees (Theorem 5.1). So, despite the negative results in the previous sections, the enumeration jump and the skip might indeed generate all natural uniformly $e$-invariant functions that are not locally constant, even if the generating process is not just iteration; we will discuss this further in the last section.
@@
If $f$ is uniformly $e$-invariant and $f(A)<_{e}A$, then $f\restriction_{\deg_{e}(A)}$ is constant. Thus, the only regressive uniformly $e$-invariant functions are locally constant.
@@
Now, let $\mathcal{T}=\{A\oplus\overline{A}\mid A\subseteq\mathbb{{N}}\}$. The function $f$ induces a uniformly $e$-invariant function $\hat{f}:\mathcal{T}\to\mathcal{{P}}(\mathbb{{N}})$ given by $\hat{f}(A\oplus\overline{A})=f(A)$. Moreover, we have that $\hat{f}\restriction_{\deg_{e}(A\oplus\overline{A})}$ is not constant for any $A\oplus\overline{A}\geq_{e}Z\oplus\overline{Z}$. It is straightforward to modify the proof of Theorem 5.1 to accept functions only defined on sets of the form $A\oplus\overline{A}$. Thus, $\hat{f}(A\oplus\overline{A})\geq_{e}A\oplus\overline{A}$ or $\hat{f}(A\oplus\overline{A})\geq_{e}(A\oplus\overline{A})^{\diamond}\equiv_{e}(A\oplus\overline{A})^{\prime}$. In any case, $A\leq_{e}A\oplus\overline{A}\leq_{e}\hat{f}(A\oplus\overline{A})=f(A)$. ∎